1. An organism is described as 2n = 58.

How many chromosomes are in one meiotic cell during telophase 1?

2. An organism is described as 2n = 58.

How many chromatids per chromosome are in one mitotic cell during metaphase?

3. An organism is described as 2n = 58.

How many chromatids per chromosome are in one meiotic cell during Anaphase 2?

4. An organism is described as 2n = 58.

How many chromatids per chromosome are in one meiotic cell during Anaphase 1?
Answer #1

Q1-

Meiosis 1 is called reductional division so the number of chromosomes will be halved in anaphase 1. so in telophase.....

58/2= 29

Q2-

A chromosome has two chromatids so in metaphase number of chromosomes would be 58*2= 116

Q3-

Since in anaphase 2 sister chromatids get separated, so number of chromatids would be 116/2= 58

Q4-

In prophase 1 the chromatids woul be double of chromosomes i.e 58*2= 96